You want to exchange 10,000 euros for Russian rubles as your travel from Paris, France to St. Petersburg, Russia.  The spot rate on the dollar/euro cross rate is $1.1878/€ and the spot rate on the ruble/dollar cross rate is Rbl 57.83/$.  How many rubles will you obtain for your euros?

Group of answer choices

205,000 Rbl.

145,580 Rbl.

486,866 Rbl.

686,905 Rbl.
Answer #1
Solution:
Answer is 4th option 686,905 Rbl.
Working Notes:
Today spot rate we have available
spot rate on the dollar/euro cross rate is $1.1878/€
Spot rate on the ruble/dollar cross rate is Rbl 57.83/$
Here We use these two rates to get cross rate Ruble per Euro
Spot rate on the dollar/euro cross rate is $1.1878/€
€1 = $1.1878
Using Unitary method we get $1 value in Euro
From 1st Equation
€1 = $1.1878
$1.1878 = €1
$1 = € (1/1.1878) 1st
Now use 2nd
Spot rate on the ruble/dollar cross rate is Rbl 57.83/$
Means $1 = Rbl 57.83 2nd
From 1st & 2nd Equation we get, as both equation $1 , must be equal
$1 = € (1/1.1878)
$1 = Rbl 57.83
From 1st & 2nd
€ (1/1.1878) = Rbl 57.83
Now we €1 value €1 = Rbl (57.83 x 1.1878)
€1 = Rbl 68.690474
So We get from 10,000 Euros from €1 = Rbl 68.690474
€1 = Rbl 68.690474
€10,000 = Rbl (68.690474 x 10,000)
€10,000 = Rbl 686904.74
€10,000 = Rbl 686,905
686,905 Rubles will be received for 10,000 Euros
Please feel free to ask if anything about above solution in comment section of the question.